OpenStudy (anonymous):

I need help with dividing complex numbers when a radical is in the numerator: \[\frac{ 2+√3i }{ 5-4i }\] I know that you multiply by conjugate of the denominator (5+4i), and you use the foil method above and below. However, the introdcution of radical 3i is what is confusing me when i combine like terms. I know the denominator will come out to 41. I started adding terms like this (below) and it all went downhill from there :) \frac{ 10+8i+5sqrt{3}i+4sqrt{3}i ^{2}}{41}

OpenStudy (anonymous):

(2+√3i )( 5 +4i ) The trouble is multiplication of these two ?

OpenStudy (anonymous):

yes that is my trouble the numerator

OpenStudy (anonymous):

Ok, so let's go as normal and see what happens.... 10 + ?

OpenStudy (anonymous):

10+8i?

OpenStudy (anonymous):

Ah, OK.....no, it is sqrt 3 not 3. -> 10 + 5sqrt 3 i + ....?

OpenStudy (anonymous):

oh i see, ok so

OpenStudy (anonymous):

AH, I see, you FOILED the other way, sorry....

OpenStudy (anonymous):

10 + 5sqrt 3 i + 8i +...?

OpenStudy (anonymous):

\[10+5\sqrt{3}i+8i+4\sqrt{3}i ^{2}\]

OpenStudy (anonymous):

what happened? Im doing something wrong?

OpenStudy (anonymous):

Yes, that's right...

OpenStudy (anonymous):

I am confusing myself, lol (I usually avoid computations)

OpenStudy (anonymous):

ok so now I need to get rid of the I and the i squared, correct?

OpenStudy (anonymous):

or is there something else I need to do first?

OpenStudy (anonymous):

Just get rid of the i^2 and bracket the i terms...

OpenStudy (anonymous):

You know what I mean, right?

OpenStudy (anonymous):

In general, you want to end up with something in the form a +ib

OpenStudy (anonymous):

\[(10+5\sqrt{3}i) (8i+4\sqrt{3}(-1))\]

OpenStudy (anonymous):

(10-4sqrt 3) + i(8+ 5 sqrt 3)

OpenStudy (anonymous):

how did you get 10-4sqrt3?

OpenStudy (anonymous):

are you foiling again?

OpenStudy (anonymous):

No, I am just arranging the terms into the form a + i b

OpenStudy (anonymous):

ok so hold

OpenStudy (anonymous):

so if i came up with \[10+8i+5\sqrt{3}i+4\sqrt{3}i ^{2}\]

OpenStudy (anonymous):

that's correct, just rearrange now...

OpenStudy (anonymous):

how would I know which constant to attach too what radical, could my asnwer be differnt

OpenStudy (anonymous):

4 sqrt 3 i^2 = -4 sqrt 3, a number. So it goes with 10, also a number.

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

and with 5rad 3 I, I can just put the I to the side since its not under the radical sign?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

ur welcome.